# Positive maps and extendibility hierarchies from copositive matrices

### v1

@@ 2.2. Convex properties of matrices
The complete characterization of the extremal rays for copositive matrices is still an open question, and the complete solution has only been found up to dimension $6$. [AHD21].
@@
as a proper cone. Indeed, according to [GB02], $I_{n^{2}}+X^{\mathsf{(C)LDUI}}_{{A},{B}}=X^{\mathsf{(C)LDUI}}_{{J_{n}+A},{I_{n}+B}}$ is separable whenever $\|(A,B)\|_{2}^{2}:=\|A\|_{2}^{2}+\|\mathring{B}\|_{2}^{2}\leq 1$, which shows that $(J_{n},I_{n})\in{\rm int}(\operatorname{\mathsf{PCP}}_{n})$. Moreover, it is obvious that $\operatorname{\mathsf{PCP}}_{n}$ cannot contain any line, proving that it is a proper cone.
@@ 2.3. Convex properties of linear maps
    A linear map $\mathcal{E}\in\mathcal{T}_{n}(\mathbb{C})$ is called completely copositive if $\mathcal{E}\circ\top$ is completely positive, or $\mathcal{E}=\mathcal{X}\circ\top$ where $\mathcal{X}$ is completely positive.
@@
Lemma 2.7 (Choi-Jamiołkowski Isomorphism [Cho75]).
@@ 2.4. Linear maps and matrices with diagonal unitary symmetry
    $\mathcal{E}(UXU^{*})=U\mathcal{E}(X)U^{*}\quad\forall X\in\mathcal{M}_{n}(\mathbb{C})\text{ and }\forall U\in D\mathcal{U}_{n},$
@@
    $\mathcal{E}(UXU^{*})=\overline{U}\mathcal{E}(X)U^{\top}\quad\forall X\in\mathcal{M}_{n}(\mathbb{C})\text{ and }\forall U\in D\mathcal{U}_{n},$
These classes of linear maps were first introduced in [Liu15, LS15]. In [JM19, SN21] , the structure of these channels was characterized completely in terms of matrix pairs, with common diagonals. Let $\mathcal{M}_{n}(\mathbb{C})^{\times 2}_{\mathbb{C}^{n}}:=\{(A,B)\in\mathcal{M}_{d}(\mathbb{C})^{\times 2}\mid\operatorname{diag}(A)=\operatorname{diag}(B)\}$
@@
    $(U\otimes U)X(U\otimes U)^{*}=X\quad\forall U\in D\mathcal{U}_{n},$
@@
    $(U\otimes\overline{U})X(U\otimes\overline{U})^{*}=X\quad\forall U\in D\mathcal{U}_{n}.$
We denote the linear space of these matrices as $\operatorname{\mathsf{LDUI}}_{n}$ and $\operatorname{\mathsf{CLDUI}}_{n}$ respectively, by adopting the notion from [SN21].
Proposition 2.12.
Define the linear bijection $J:\mathcal{T}_{n}(\mathbb{C})\rightarrow\mathcal{M}_{n}(\mathbb{C})\otimes\mathcal{M}_{n}(\mathbb{C})$ as $J(\Phi)=\sum_{i,j=1}^{n}\Phi(|i\rangle\langle j|)\otimes|i\rangle\langle j|$. Then, $\Phi\in\mathcal{T}_{n}(\mathbb{C})$
  • is DUC if and only if $J(\Phi)$ is CLDUI. In particular, we have
@@
  • is CDUC if and only if $J(\Phi)$ is LDUI. In particular, we have
@@
Consider the following cone:
@@
$(A,B)\in\operatorname{\mathsf{CLDUI}}_{n}^{+}\iff\Phi^{\mathsf{DUC}}_{{A},{B}}\text{ is completely positive}\iff\Phi^{\mathsf{CDUC}}_{{A},{B}}\text{ is completely copositive}.$
@@
The inclusion follows from easily from Definition 3.1, while the equivalence is proven in [SN21, Lemma 2.12], see also [NS21, Lemma 7.6] or [JM19, Theorem 5.2]
@@
Consider the following cone:
@@
$(A,B)\in\mathsf{PDNN}_{n}\iff\Phi^{\mathsf{(C)DUC}}_{{A},{B}}\text{ is PPT}.$
@@
Bipartite Hermitian matrices which are diagonal in the Dicke basis
$X=\sum_{1\leq i\leq j\leq n}Y_{ij}|\psi_{ij}\rangle\langle\psi_{ij}|\in\mathcal{B}(\mathbb{C}^{n}\vee\mathbb{C}^{n}),\quad Y_{ij}\in\mathbb{R},$
@@ 3. Pairwise copositivity
$\langle vw|X^{\mathsf{CLDUI}}_{{A},{B}}|vw\rangle$ $=\langle v\odot\overline{v}|A|w\odot\overline{w}\rangle+\langle v\odot w|\mathring{B}|v\odot w\rangle.$ (9)
@@
We show the proof of the first equation by the following computations.
@@
Here, $\mathcal{T}_{\operatorname{\mathsf{LDUI}}}(X)=\mathcal{T}_{D\mathcal{U}_{n}^{\otimes 2}}(X):=\int_{D\mathcal{U}_{n}}(U\otimes U)X(U\otimes U)^{*}$ is the LDUI-twirling, and we use Eq. 7 to compute the inner product (We refer to [SN21, PJPY24] for general properties of twirling maps).
The second formula can be shown similarly. ∎
@@
$(A,B)\in\operatorname{\mathsf{PCOP}}_{n}\iff\Phi^{\mathsf{(C)DUC}}_{{A},{B}}\text{ is positive}$
@@
Assume that $(A,B)\in\operatorname{\mathsf{PCOP}}_{n}$. Then, by the Proposition 3.3, we get
@@
This is equivalent to the condition that that $X^{\mathsf{LDUI}}_{{A},{B}}$ is block positive $\iff\Phi^{\mathsf{CDUC}}_{{A},{B}}$ is positive [EK00]. ∎
We collect in the following proposition some simple results about the convex cone $\operatorname{\mathsf{PCOP}}$.
@@
[…]
$\braket{i|A|j}+0=A_{ij}\geq 0.$
For Item 2, we replace in Eq. 10, $v=z,w=\bar{z}\in\mathbb{C}^{n}$.We get
@@
  • Item 4: A “classical” map $\Phi^{\mathsf{(C)DUC}}_{{A},{\operatorname{diag}(A)}}(X)=\operatorname{diag}\{A\operatorname{diag}[X]\}$ is positive $\iff$ it is completely positive $\iff$ it is entanglement breaking.
  • Item 5: A (transpose-)Schur product map $\Phi^{\mathsf{DUC}}_{{\operatorname{diag}(B)},{B}}(X)=B\odot X$ (resp. $\Phi^{\mathsf{CDUC}}_{{\operatorname{diag}(B)},{B}}(X)=B\odot X^{\top}$) is positive $\iff$ it is completely positive (resp. completely co-positive).
  • Item 6: A (transpose-)Schur product map $\Phi^{\mathsf{CDUC}}_{{\operatorname{diag}(B)},{B}}(X)=B\odot X^{\top}$ is entanglement breaking $\iff$ it is “classical” (i.e. only takes into account the diagonal of the input).
@@
We have seen in Proposition 2.1 that $(A,A)\in\operatorname{\mathsf{PCP}}_{n}\iff A\in\mathsf{CP}_{n}$. From Item 7 above, it follows that a similar property does not hold for $\operatorname{\mathsf{PCOP}}_{n}$ and $\operatorname{\mathsf{COP}}_{n}$. In the language of linear maps, the result from Item 7 translates to the fact that the linear $\mathsf{DUC}$ map
@@
In particular, $(\mathring{J},-\mathring{J})\in\operatorname{\mathsf{PCOP}}_{n}$. Note that since $\mathring{N}\notin\mathsf{PSD}^{\mathbb{C}}_{n}$ unless $\mathring{N}=0$, we have

### v4

@@ 2.2. Convex properties of matrices
The complete characterization of the extremal rays for copositive matrices is still an open question, and the complete solution has only been found up to dimension $6$ [Hil12, AHD21].
@@
as a proper cone. Indeed, we will see later that $(J_{n}+A,I_{n}+B)\in\operatorname{\mathsf{PCP}}_{n}$ whenever
$\|(A,B)\|_{2}^{2}:=\|A\|_{2}^{2}+\|\mathring{B}\|_{2}^{2}=\sum_{i,j}|A_{ij}|^{2}+\sum_{i\neq j}|B_{ij}|^{2}\leq 1,$
which implies that $(J_{n},I_{n})\in{\rm int}(\operatorname{\mathsf{PCP}}_{n})$, from the correspondence with the separability of diagonal unitary invariant states (Proposition 2.13) and the separability criterion in [GB02, Theorem 1]. Moreover, it is obvious that $\operatorname{\mathsf{PCP}}_{n}$ cannot contain any line, proving that it is a proper cone.
@@ 2.3. Convex properties of linear maps
    A linear map $\mathcal{E}\in\mathcal{T}_{n}(\mathbb{C})$ is called completely copositive if $\mathcal{E}\circ\top$ is completely positive, or equivalently, $\mathcal{E}=\mathcal{X}\circ\top$ where $\mathcal{X}$ is completely positive.
@@
Lemma 2.7 (Choi-Jamiołkowski Isomorphism [Cho75a]).
@@ 2.4. Linear maps and matrices with diagonal unitary symmetry
    $\forall X\in\mathcal{M}_{n}(\mathbb{C})\text{ and }\forall U\in D\mathcal{U}_{n},\quad\mathcal{E}(UXU^{*})=U\mathcal{E}(X)U^{*}$
@@
    $\forall X\in\mathcal{M}_{n}(\mathbb{C})\text{ and }\forall U\in D\mathcal{U}_{n},\quad\mathcal{E}(UXU^{*})=\overline{U}\mathcal{E}(X)U^{\top}$
These classes of linear maps were first introduced in [Liu15, LS15]. In [JM19, SN21] , the structure of these channels was characterized completely in terms of matrix pairs, with common diagonals. Let $\mathcal{M}_{n}(\mathbb{C})^{\times 2}_{\mathbb{C}^{n}}:=\{(A,B)\in\mathcal{M}_{d}(\mathbb{C})^{\times 2}\mid\operatorname{diag}(A)=\operatorname{diag}(B)\}$ denote the set of such matrix pairs.
@@
    $\forall U\in D\mathcal{U}_{n},\quad(U\otimes U)X(U\otimes U)^{*}=X.$
@@
    $\forall U\in D\mathcal{U}_{n},\quad(U\otimes\overline{U})X(U\otimes\overline{U})^{*}=X.$
We denote these linear spaces by $\operatorname{\mathsf{LDUI}}_{n}$ and $\operatorname{\mathsf{CLDUI}}_{n}$, following the terminology of [SN21]. We then further identify $\mathcal{M}_{n}(\mathbb{C})^{\times 2}_{\mathbb{C}^{n}}\cong\operatorname{\mathsf{LDUI}}_{n}\cong\operatorname{\mathsf{CLDUI}}_{n}$ via the Choi-Jamiołkowski isomorphism (Lemma 2.7) described below (see also Proposition 2.14 for related adapted notation).
Proposition 2.12 ([SN21]).
For $\Phi\in\mathcal{T}_{n}(\mathbb{C})$,
  • $\Phi$ is DUC if and only if $J(\Phi)$ is CLDUI. In particular, we have
@@
  • $\Phi$ is CDUC if and only if $J(\Phi)$ is LDUI. In particular, we have
@@
Consider the following cones:
$\operatorname{\mathsf{LDUI}}^{+}_{n}$ $:=\Big\{(A,B)\in\mathcal{M}_{n}(\mathbb{R})\underset{\mathbb{R}^{n}}{\times}\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{C})\,:\,A\in\operatorname{\mathsf{EWP}}_{n},\,A_{ij}A_{ji}\geq|B_{ij}|^{2}\,\forall\,i\neq j\,\Big\}.$
@@
$(A,B)\in\operatorname{\mathsf{LDUI}}_{n}^{+}$ $\iff\Phi^{\mathsf{CDUC}}_{{A},{B}}\text{ is completely positive}$
$\iff\Phi^{\mathsf{DUC}}_{{A},{B}}\text{ is completely copositive}\iff X^{\mathsf{LDUI}}_{{A},{B}}\geq 0,$
$(A,B)\in\operatorname{\mathsf{CLDUI}}_{n}^{+}$ $\iff\Phi^{\mathsf{DUC}}_{{A},{B}}\text{ is completely positive}$
$\iff\Phi^{\mathsf{CDUC}}_{{A},{B}}\text{ is completely copositive}\iff X^{\mathsf{CLDUI}}_{{A},{B}}\geq 0.$
@@
The inclusions follows from easily from Definition 3.1, while the equivalence is proven in [SN21, Lemma 2.12], see also [NS21, Lemma 7.6] or [JM19, Theorem 5.2]
@@
Consider the following cone of pairwise doubly non-negative matrices:
@@
$(A,B)\in\mathsf{PDNN}_{n}\iff\Phi^{\mathsf{(C)DUC}}_{{A},{B}}\text{ is PPT}\iff X^{\mathsf{(C)LDUI}}_{{A},{B}}\text{ is PPT}.$
@@
Bipartite positive semidefinite matrices which are diagonal in the Dicke basis
$X=\sum_{1\leq i\leq j\leq n}Y_{ij}|\psi_{ij}\rangle\langle\psi_{ij}|\in\mathcal{B}(\mathbb{C}^{n}\vee\mathbb{C}^{n}),\quad Y_{ij}\in\mathbb{R}_{+},$
@@ 3. Pairwise copositivity
$\langle vw|X^{\mathsf{CLDUI}}_{{A},{B}}|vw\rangle$ $=\langle v\odot\overline{v}|A|w\odot\overline{w}\rangle+\langle v\odot w|\mathring{B}|v\odot w\rangle,$ (9)
for $v,w\in\mathbb{C}^{n}$ and $|vw\rangle:=|v\otimes w\rangle$.
@@
We first consider the LDUI-twirling
$\mathcal{T}_{\operatorname{\mathsf{LDUI}}}(X)=\mathcal{T}_{D\mathcal{U}_{n}^{\otimes 2}}(X):=\int_{D\mathcal{U}_{n}}(U\otimes U)X(U\otimes U)^{*}dU,\quad X\in\mathcal{M}_{n}(\mathbb{C})^{\otimes 2},$ (10)
where $dU$ denotes the normalized Haar measure on $D\mathcal{U}_{n}$. Then $\mathcal{T}_{\operatorname{\mathsf{LDUI}}}$ can be considered as a orthogonal projection of $\mathcal{M}_{n}(\mathbb{C})^{\otimes 2}$ onto the linear space $\operatorname{\mathsf{LDUI}}_{n}$ (we refer to [SN21, PJPY24] for general properties of twirling maps). Furthermore, we have
$\forall\,X\in\mathcal{M}_{n}(\mathbb{C})^{\otimes 2},\quad$ $X\in\operatorname{\mathsf{LDUI}}_{n}\iff\mathcal{T}_{\operatorname{\mathsf{LDUI}}}(X)=X,$
[…]
$\forall\,v,w\in\mathbb{C}^{n},\quad$ $\mathcal{T}_{\operatorname{\mathsf{LDUI}}}(|vw\rangle\langle vw|)=X^{\mathsf{LDUI}}_{{|v\odot\bar{v}\rangle\langle w\odot\bar{w}|},{|v\odot\bar{w}\rangle\langle v\odot\bar{w}|}},$
where the first two facts follows from [PJPY24, Section 2.3] and the last fact is from [SN21, Remark 5.16]. Now we show the proof of the first equation by the following computations.
@@
The second formula can be shown similarly via the CLDUI-twirling
$\mathcal{T}_{\mathsf{CLDUI}}(X):=\int_{D\mathcal{U}_{n}}(U\otimes\overline{U})X(U\otimes\overline{U})^{*}dU,\quad X\in\mathcal{M}_{n}(\mathbb{C})^{\otimes 2},$ (11)
which is left to the reader. ∎
@@
$(A,B)\in\operatorname{\mathsf{PCOP}}_{n}\iff\Phi^{\mathsf{(C)DUC}}_{{A},{B}}\text{ is positive}\iff X^{\mathsf{(C)LDUI}}_{{A},{B}}\text{ is block-positive}.$
@@
Assume that $(A,B)\in\operatorname{\mathsf{PCOP}}_{n}$. Then, by Proposition 3.3, we get
@@
This is equivalent to the condition that that $X^{\mathsf{LDUI}}_{{A},{B}}$ is block-positive $\iff\Phi^{\mathsf{CDUC}}_{{A},{B}}$ is positive [EK00]. ∎
We further collect in the following proposition some simple results about the convex cone $\operatorname{\mathsf{PCOP}}$.
@@
    $(D_{1}^{*}D_{1}AD_{2}^{*}D_{2},D_{1}^{*}D_{2}^{*}BD_{1}D_{2})\in\operatorname{\mathsf{PCOP}}_{n}$ for any complex diagonal matrices $D_{1}$ and $D_{2}$.
@@
For Items 1 and 2, we refer to [SN21, Proposition 6.7].
To show Item 3, we replace in Eq. 12, $v=z,w=\bar{z}\in\mathbb{C}^{n}$.We get
@@
Item 4 follows from the substitution $v\mapsto D_{1}v=\operatorname{diag}[D_{1}]\odot v$ and $w\mapsto D_{2}w=\operatorname{diag}[D_{2}]\odot w$ in Eq. 12, in which case $D_{1}v\odot\overline{D_{1}v}=D_{1}^{*}D_{1}(v\odot\overline{v})$ and $D_{1}v\odot D_{2}w=D_{1}D_{2}(v\odot w)$.
@@
Finally, Item 8 follows from Item 1 and Proposition 2.14. ∎
@@
  • Item 5: A “classical” map $\Phi^{\mathsf{(C)DUC}}_{{A},{\operatorname{diag}(A)}}(X)=\operatorname{diag}\{A\operatorname{diag}[X]\}$ is positive $\iff$ it is completely positive $\iff$ it is entanglement breaking.
  • Item 6: A (transpose-)Schur product map $\Phi^{\mathsf{DUC}}_{{\operatorname{diag}(B)},{B}}(X)=B\odot X$ (resp. $\Phi^{\mathsf{CDUC}}_{{\operatorname{diag}(B)},{B}}(X)=B\odot X^{\top}$) is positive $\iff$ it is completely positive (resp. completely co-positive).
  • Item 7: A (transpose-)Schur product map $\Phi^{\mathsf{CDUC}}_{{\operatorname{diag}(B)},{B}}(X)=B\odot X^{\top}$ is entanglement breaking $\iff$ it is “classical” (i.e. only takes into account the diagonal of the input).
@@
We have seen in Proposition 2.1 that $(A,A)\in\operatorname{\mathsf{PCP}}_{n}\iff A\in\mathsf{CP}_{n}$. From Item 8 above, it follows that a similar property does not hold for $\operatorname{\mathsf{PCOP}}_{n}$ and $\operatorname{\mathsf{COP}}_{n}$. In the language of linear maps, the result from Item 8 translates to the fact that the linear $\mathsf{DUC}$ map
@@
In particular, $(\mathring{J}_{n},-\mathring{J}_{n})\in\operatorname{\mathsf{PCOP}}_{n}$. Note that since $\mathring{N}\notin\mathsf{PSD}^{\mathbb{C}}_{n}$ unless $\mathring{N}=0$, we have
@@
We complete this section by providing the characterizations of pairwise copositivity in terms of copositivity, which generalize Item 3 of Proposition 3.5. First, for $A\in\operatorname{\mathsf{EWP}}_{n}$, let us denote by
$\widetilde{A}:=(A\odot A^{\top})^{\odot 1/2}=(\sqrt{A_{ij}A_{ji}})_{1\leq i,j\leq n}.$
Theorem 3.9.
If a pair $(A,B)\in\mathcal{M}_{n}(\mathbb{R})\underset{\mathbb{R}^{n}}{\times}\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{C})$ satisfies $A\in\operatorname{\mathsf{EWP}}_{n}$ and $\widetilde{A}+{\rm Re}(U^{*}\mathring{B}U)\in\operatorname{\mathsf{COP}}_{n}$ for every $U\in D\mathcal{U}_{n}$, then $(A,B)\in\operatorname{\mathsf{PCOP}}_{n}$. In particular, for $A\in\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{R})$ real symmetric,
$(A,B)\in\operatorname{\mathsf{PCOP}}_{n}\iff A\in\operatorname{\mathsf{EWP}}_{n}\text{ and }A+{\rm Re}(U^{*}\mathring{B}U)\in\operatorname{\mathsf{COP}}_{n}\;\;\forall\,U\in D\mathcal{U}_{n}$
Proof.
For $v,w\in\mathbb{C}^{n}$ , let us write $v\odot w=Up$ where $p\in\mathbb{R}_{+}^{n}$ and $U\in D\mathcal{U}_{n}$. Then we have
$\langle v\odot\bar{v}|A\big|w\odot\bar{w}\rangle+\langle v\odot w|\mathring{B}|v\odot w\rangle$ $=\sum_{i,j}A_{ij}|v_{i}|^{2}|w_{j}|^{2}+\langle p|U^{*}\mathring{B}U|p\rangle$
$=\sum_{i<j}(A_{ij}p_{j}^{2}\frac{|v_{i}|^{2}}{|v_{j}|^{2}}+A_{ji}p_{i}^{2}\frac{|v_{j}|^{2}}{|v_{i}|^{2}})+\langle p|U^{*}BU|p\rangle$
$\geq\sum_{i<j}2\sqrt{A_{ij}A_{ji}}p_{i}p_{j}+\langle p|U^{*}BU|p\rangle$
$=\langle p\big|\widetilde{A}+{\rm Re}(U^{*}\mathring{B}U)\big|p\rangle\geq 0,$
where in the second inequality, we use that $\operatorname{diag}(A)=\operatorname{diag}(B)=U^{*}\operatorname{diag}(B)U$ and in the next step the AM-GM inequality. This shows the first assertion.
[…]
The corollary below would play a key role in the next section. Let us first denote by $D\mathcal{O}_{n}:=\big\{O\in D\mathcal{U}_{n}:\operatorname{diag}[O]\in\{\pm 1\}^{n}\big\}$ the finite group of diagonal orthogonal matrices.
Corollary 3.10.
If $(A,B)$ satisfies $A\in\operatorname{\mathsf{EWP}}_{n}$, $B\in\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{R})$, and $\widetilde{A}+O\mathring{B}O\in\operatorname{\mathsf{COP}}_{n}$ for every $O\in D\mathcal{O}_{n}$, then $(A,B)\in\operatorname{\mathsf{PCOP}}_{n}$. In particular, for $A,B\in\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{R})$ real symmetric,
$(A,B)\in\operatorname{\mathsf{PCOP}}_{n}\iff A\in\operatorname{\mathsf{EWP}}_{n}\text{ and }A+O\mathring{B}O\in\operatorname{\mathsf{COP}}_{n}\;\;\forall\,O\in D\mathcal{O}_{n}.$
Proof.
We only need to show the first assertion. First of all, for all $D$ real diagonal, we can write $D=OP=PO$ for some $P$ a nonnegative diagonal matrix and $O\in D\mathcal{O}_{n}$. Then $|D|:=\sqrt{D^{*}D}=P$ and hence the assumption implies that
$|D|\widetilde{A}|D|+D\mathring{B}D=P(\widetilde{A}+O\mathring{B}O)P\in\operatorname{\mathsf{COP}}_{n}.$ (14)
Now for $U\in D\mathcal{U}_{n}$, let us write $U=D_{1}+iD_{2}\in D\mathcal{U}_{n}$ with $D_{1},D_{2}$ real diagonal (hence $D_{1}^{2}+D_{2}^{2}=I_{n}$). We claim that for all $A\in\mathsf{EWP}_{n}$, we have $\widetilde{A}-|D_{1}|\widetilde{A}|D_{1}|-|D_{2}|\widetilde{A}|D_{2}|\in\operatorname{\mathsf{EWP}}_{n}^{\mathrm{sa}}$. To show this, we use the AM-GM inequality and the fact that $D_{1}^{2}+D_{2}^{2}=I_{n}$:
$\forall i,j\in[n],\qquad|D_{1}|_{ii}|D_{1}|_{jj}+|D_{2}|_{ii}|D_{2}|_{jj}\leq\frac{|D_{1}|_{ii}^{2}+|D_{1}|_{jj}^{2}}{2}+\frac{|D_{2}|_{ii}^{2}+|D_{2}|_{jj}^{2}}{2}=1.$ (15)
Then, since $B$ is real, we have
$\widetilde{A}+{\rm Re}(U^{*}\mathring{B}U)$ $=\widetilde{A}+D_{1}\mathring{B}D_{1}+D_{2}\mathring{B}D_{2}$
[…]
This shows that $\widetilde{A}+{\rm Re}(U^{*}\mathring{B}U)$ is copositive, and then the first claim follows now from Theorem 3.9. ∎